Consider the following scenario:
• Let P(C) = 0.4
• Let P(D) = 0.5
• Let P(C|D) = 0.6
Find P(C AND D).

Answer by ikleyn(52876)

.

By the definition of the CONDITIONAL PROBABILITY,


    P(C|D) = P(C AND D) / P(D).


So, we have  0.6 = P(C AND D) / 0.5.


It implies  P(C AND D) = 0.6*0.5 = 0.3.    ANSWER

Solved.

Notice that this info  " Let P(C) = 0.4 "  is excessive;  it was not used in the solution,  and therefore is  IRRELEVANT  to the problem.